Use fractions to determine which wrench is larger.
3/8 0r 7/16?
Found 2 solutions by RAY100, rapaljer: Answer by RAY100(1637) 
 In fractions the numerator is the number of pieces,,,the denominator is the size of the pieces.
.
To compare, usually the denominator is the same
.
3/8 vs 7/16,,,,,lets use LCD = 16
.
3/8 = 6/16....16/8=2,,,2*3=6
.
now problem is 3/8 =6/16 vs 7/16,,,,,with 7/16 Larger than 6/16 or 3/8
.
7/16 > 3/8
Answer by rapaljer(4671) 
 You can either convert these to decimal form, which allows you to compare and see which is larger. Do you have a calculator? If so, this is the easiest way. Just put 3 divided by 8 in the calculator (it gives you .375) and then do 7 divided by 16 (this should give you .4375). Of course 7/16 is larger.
As a second method, you can find a common denominator, which in this case is 16. Then multiply the numerator and denominator by 2, so the denominators are the same. This gives you 3/8= 6/16. You can see that 7/16 is more than 6/16.
R^2
Dr. Robert J. Rapalje
| |
|